OpenStudy (anonymous):

Find the sum of the arithmetic sequence. 17, 19, 21, 23, ..., 35 PLEASE

OpenStudy (anonymous):

Answer is 260 (: I can't explain, i suck at explaining dude, sorry. plus i gotta go.

OpenStudy (anonymous):

Hope I helped you.

OpenStudy (anonymous):

gives the derivation

OpenStudy (jdoe0001):

so we use the equation to find a sequence term as => \(\bf \Large a_n = a_1+(n+1)d\)

OpenStudy (jdoe0001):

so looking at -> 17, 19, 21, 23, ..., 35 <- we see the 1st term, \(\bf a_1 = 17\) the common ratio is "2", 2 is being added to the previous term to get the next we also see a 35 in the end, what term is that? dunno but we know that \(\bf a_n = a_1+(n+1)d\) so let's use those values and solve for "n" so we can see what that term is \(\bf a_n = a_1+(n+1)d \implies 35 = 17+(n+1)2\\ 35 = 17+ 2n+2 \implies 35 = 2n+19 \implies 16 = 2n\\ 8 = n \textit{, so we now know that 35 is the } 8^{th}\ term\)

OpenStudy (jdoe0001):

so you'd think, who cares? well, now we'll use the Sum of a partial sequence equation of \(\bf \Large S_n = \cfrac{n}{2}(a_1+a_n)\)

OpenStudy (jdoe0001):

so we know that our end term in the sequence is the 8th term, 35 we know our 1st term, 17 thus \(\bf S_n = \cfrac{n}{2}(a_1+a_n) \implies S_8 = \cfrac{8}{2}(17+35)\)

OpenStudy (anonymous):

Ahhh okay! i am understanding, I had no idea how to get the n :D

OpenStudy (anonymous):

S8 would be 208?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

what do we need s8 for?

OpenStudy (jdoe0001):

s8 is the "sum of the arithmetic sequence"

OpenStudy (jdoe0001):

it's just the notation saying, Sum of 8 terms of this sequence

OpenStudy (anonymous):

ahhh but that isn't the whole sum right? bc that's not an answer choice

OpenStudy (jdoe0001):

well, that's what I got from the Sum equation.... so what choices do you have?

OpenStudy (anonymous):

260 179 37 160